OpenStudy (anonymous):

Let y = ((ln(x^2-1)^-3)/2). Find the derivative

OpenStudy (anonymous):

\[\frac{ \ln(x ^{2}-1)^{-3} }{ 2 }\] \[\frac{ 1 }{ 2 }(-3)\ln(x ^{2}-1)^{-4}(\frac{ 1 }{ x ^{2}-1 })(2x)\] \[\frac{ -3x }{ \ln(x ^{2}-1)^{4} (x ^{2}-1)}\]

OpenStudy (anonymous):